OpenStudy (anonymous):

Suppose p varies directly with d, and p = 3 when d = 5. What is the value of d when p = 12?

OpenStudy (anonymous):

@tester97

OpenStudy (tester97):

I'm not sure this is correct but the way i got that was to set up an equation and it looked like this \[d=\frac{ ? }{ p }\] then i put in the actual numbers and it looked like this \[5=\frac{ ? }{ 3 }\] and to find the numerator i just multiplied 5 and 3 and it looked like this \[5=\frac{ 15 }{ 3 }\] so then i changed out the 3 and put 12 and it looked like this \[?=\frac{ 15 }{ 12 }\] and you just divide that and you get \[1.25\]